# Dielectronic recombination studies on Fe2+

### v1

@@ I Introduction
$3d^{6}~{}(i)+e^{-}\to 3d^{5}4s4d~{}(a)~{}\&~{}3d^{5}4p^{2}~{}(a)$
$\to 3d^{5}4s4p~{}(d)~{}\&~{}3d^{6}4p~{}(d)$ (1)
$\to 3d^{6}4s~{}(f)$
Here, i denotes the ground state of Fe III, whereas a and d represent the autoionization and the radiative states of Fe II, and f denotes the final state (ground state) of Fe II. Fig. 1 provides a schematic diagram of the various stages of the DR process considered in the present work.
@@ II Theory and calculation of resonance strengths
As for the radial wave functions, the following set of MCDHF equations can be derived [47]:
$\left(\frac{d}{dr}+\frac{\kappa_{a}}{r}\right)P_{n_{a}\kappa_{a}}(r)-\left(2c-\frac{\epsilon_{a}}{c}+\frac{Y_{a}(r)}{cr}\right)Q_{n_{a}\kappa_{a}}(r)$
$=-\frac{\chi_{a}^{P}(r)}{r},\quad$ (12a)
$\left(\frac{d}{dr}-\frac{\kappa_{a}}{r}\right)Q_{n_{a}\kappa_{a}}(r)+\left(-\frac{\epsilon_{a}}{c}+\frac{Y_{a}(r)}{cr}\right)P_{n_{a}\kappa_{a}}(r)$
$=\frac{\chi_{a}^{Q}(r)}{r}\,.\quad$ (12b)
In this equation, $a$ is the index for a given orbital. The equations are coupled via the direct ($Y_{a}$) and exchange ($\chi_{a}^{P},\chi_{a}^{Q}$) potentials [47]. The equations (12) and (12) are solved iteratively together with the algebraic equations (11) until self-consistency is reached.
@@
The initial state of the DR process which consists of the ground-state ion and a continuum electron with an asymptotic momentum $\vec{p}$ and spin projection $m_{s}$. In addition, $\Gamma_{d}$ is the total natural width of the intermediate autoionizing state, which is the sum of the radiative and autoionization widths: $\Gamma_{d}=A^{d}_{r}+A^{d}_{a}$ (note that rates and corresponding line widths are equivalent in atomic units). $L_{d}(E)$ is the Lorentzian line shape function, expressed as
@@
The total rate coefficients for thermal plasmas are described by the following equation, derived by summing across all possible autoionization channels and averaging over the Maxwellian distribution of electron energies [55]. In this expression, $k$ denotes the Boltzmann constant, $h$ is the Planck constant, $T$ represents the electron temperature, and $E$ is the resonance energy:
@@ III Results
The MCDHF method, combined with the relativistic configuration interaction (RCI) approach, is used to calculate the energies of the auto-ionizing and radiative decay states of Fe+, as well as the ground states of Fe2+, as shown in Table 1 and Table 2, respectively. The RCI calculations account for the Breit interaction and approximate quantum electrodynamic (QED) terms using the rci code [39].
Transition Present Nahar [56] Johansson ([57]
and ref. therein)
$3d^{6}4s$[6D] $\rightarrow$ $3d^{6}$[5D] 15.9171 16.0302 16.1353
$3d^{6}$[5D] $\rightarrow$ $3d^{5}4s4d$[6D] 1.8298 1.9674 1.7633
$3d^{6}$[5D] $\rightarrow$ $3d^{5}4p^{2}$[6S] 1.7069
$3d^{5}4s4p$[8P] $\rightarrow$ $3d^{6}$[5D] 9.9064 10.5893 9.6562
$3d^{6}4p$[4F] $\rightarrow$ $3d^{6}$[5D] 10.5517 10.8410 10.6437
@@
Since electron correlations play a crucial role in this system, the ASFs are expanded using a large CSF basis. In the present calculation scheme, the CSFs are generated via single and double (SD) excitations of electrons from the Fe2+ ($3d^{6}$ [5D]) valence orbitals to auto-ionizing states of Fe+ ($3d^{5}4s4d$ [6D] and $3d^{5}4p^{2}$ [6S]). A similar approach is applied for radiative decay states, where SD excitations of electrons from the $3d^{6}$ [5D] valence orbitals to the radiative decay states $3d^{5}4s4p$ [8P] and $3d^{6}4p$ [4F] are considered. This method is also used to calculate transition energies from the ground state of Fe+ ($3d^{6}4s$ [6D]) to Fe2+ ($3d^{6}$ [5D]). In these calculations, the spectroscopic orbitals in $3d$, $4s$, and $4p$ were kept open, depending on the states under consideration. Valence-valence and core-valence correlations were included, and the convergence of the transition energies and radiative decay rates was monitored through a layer-by-layer expansion. The active space of virtual orbitals was optimized up to the $8k$ subshell. An example of this convergence is shown in Figure 3, where the energies of the $3d^{5}4s4p$ [8P] and $3d^{6}4p$ [4F] states are seen to stabilize around the virtual orbital principal quantum number $n=8$. We compare our result with the theoretical work of Nahar and Pradhan [56] and experimental work of S. Johansson ([57] and references therein), and it is observed that the present values agree well with the experimental value of S. Johansson ([57] and references therein) and demonstrate an improvement upon earlier theoretical work [56].
For the $3d^{6}$ [5D] state with total angular momentum $J=4$, the total number of CSFs considered was 279,863. For the states $3d^{6}4s$ [6D] ($J=1/2$ to $9/2$), $3d^{5}4s4d$ [6D] ($J=1/2$ to $9/2$), $3d^{5}4p^{2}$ [6S] ($J=5/2$), $3d^{5}4s4p$ [8P] ($J=1/2$ to $17/2$), and $3d^{6}4p$ [4F] ($J=1/2$ to $15/2$), the total number of CSFs included were 987,905; 5,714,055; 1,034,395; 5,986,145; and 2,955,693, respectively.
Config. state Present (eV) Ekberg (NIST) [58] (eV)
@@
Figure 3: Calculated energies (in eV) for the states $3d^{5}4s4p$[8P] and $3d^{6}4p$[4F] monitored via layer-by-layer expansion. Black and red lines indicate present and previous studies done on the $3d^{5}4s4p$[8P] and $3d^{6}4p$[4F] states respectively.
@@
The ionization energy of Fe+ ($3d^{6}4s$ [6D]) to Fe2+ ($3d^{6}$ [5D]) is calculated to be 15.917eV, differing by only 1.3% from the experimental values (see[57] and references therein). The autoionization level energies of the states $3d^{5}4s4d$ [6D] and $3d^{5}4p^{2}$ [6S] from $3d^{6}$ [5D] are calculated to be 1.830eV and 1.707eV, respectively, agreeing within 3.7% of the experimental energies. Similarly, our calculated radiative decay levels of $3d^{5}4s4p$ [8P] and $3d^{6}4p$ [4F] from $3d^{6}$ [5D] are 9.9064eV and 10.5517eV, respectively, differing by 2.5% and 0.8% from experimental values. Since our calculations incorporate a more comprehensive set of relevant CSFs, the present results are expected to be more accurate than previous theoretical predictions [56]. These energy levels for various DR states are graphically represented in Fig.2. Table 1 provides a detailed comparison of these energy values with earlier studies. It is evident from the table that, apart from the $3d^{6}4s$ [6D] $\rightarrow$ $3d^{6}$ [5D] transitions, the present results show better agreement with experimental values compared to prior theoretical data.
@@
The ground-state energies for Fe2+ are also calculated and presented in Table 2. These values are compared with the experimental data provided by Ekberg [58] in the NIST database. The calculated results exhibit excellent agreement with experimental data across all states. As discussed later, these energies were subsequently used in the FAC to generate accurate DR spectra. This underscores the enhanced accuracy and reliability of the current computational approach, demonstrating its capability to predict transitions and align well with experimental observations.
@@
Figure 4: Total DR cross section for Fe2+ recombining into Fe+ is plotted against the relative electron energy. The Lorentzian line shapes (eq.  14) are convoluted with a Gaussian function with a 10-meV width for a better representation of the spectra. The inset presents an enlarged view of the spectrum in the range from 0 to 0.25 eV.
@@
We present the total DR cross section for Fe2+ recombining to Fe+, illustrated in Figure 4. The calculations employ the FAC code to derive the DR spectrum for two autoionizing states of Fe+, namely, $3d^{5}4s4d$ and $3d^{5}4p^{2}$. To enhance the resolution, Lorentzian line shapes, as described in equation 14, are convolved with a Gaussian function characterized by a width of 10 meV. The resulting spectrum exhibits pronounced resonances below 1 eV for both autoionizing states (especially for the $3d^{5}4s4d$ state), with a gradual attenuation of the resonances. Notably, the contributions from the $3d^{5}4s4d$ state are much stronger than from the $3d^{5}4p^{2}$ state. We observe a significant resonance peak at around 0.05 eV, attributed to the $3d^{5}4s4d$ state, along with a couple of resonance peaks between 0.1 to 0.2 eV. Beyond 0.2 eV, the spectrum is almost completely attenuated, with a small resonance peak at 0.4 eV due to the $3d^{5}4p^{2}$ state. There is also a small resonance peak at 0.66 eV, resulting from the contribution of both autoionizing states $3d^{5}4s4d$ and $3d^{5}4p^{2}$. Since the energy levels calculated by FAC are not very accurate and have an esimated uncertainty of a few eV, these energies require some adjustments. Therefore, we adjusted these energies to the more accurate energies calculated using the large-scale MCDHF method as implemented by the GRASP code. These adjustments were necessary to obtain more reliable energy positions for low-lying DR resonances. The use of the MCDHF method in calculating resonance energies and radiative rates has proven to be effective in the past, yielding results that align well with experimental data [59, 60, 61]. This agreement underscores the validity of the theoretical approach employed in this study. Since a significant amount of data was generated in the present calculation, the energies, resonance strengths, along with their respective widths, are provided as supplementary data.
@@
The total DR rate coefficients for Fe2+ are shown in the figure as a function of temperature on Fig. 5. At high temperatures, the DR rate coefficient decreases according to a $\sim T^{-3/2}$ dependence. At low temperatures, however, the DR process is dominated by a few resonances located just above the ionization limit, which contribute significantly to the rate at low temperatures. This occurs because the $\sim T^{-3/2}$ factor compensates for the low temperature. However, as the temperature gets very low, the exponential cut-off eventually overwhelms the $\sim T^{-3/2}$ increase, causing the rate to decline sharply.

### v2

@@ I Introduction
$3d^{6}~{}(i)+e^{-}\to 3d^{5}4s4d~{}(d)~{}\&~{}3d^{5}4p^{2}~{}(d)$
$\to 3d^{5}4s4p~{}(f)~{}\&~{}3d^{6}4p~{}(f)$ (1)
$\to 3d^{6}4s~{}(f^{\prime})$
Here, i denotes the initial state, which corresponds to the ground state of Fe III. The labels d and f represent the autoionizing state of Fe II and the final bound state of Fe II (with no further autoionization involved), respectively. Additionally, $f^{\prime}$ denotes the ground state of Fe II. Figure 1 provides a schematic representation of the various stages of the DR process considered in this work.
@@ II Theory and calculation of resonance strengths
As for the radial wave functions, the set of MCDHF equations given in  [45] are solved iteratively together with the algebraic equations (11) until self-consistency is reached.
@@
The initial state of the DR process consists of the ground-state ion and a continuum electron with an asymptotic momentum $\vec{p}$ and spin projection $m_{s}$. In addition, $\Gamma_{d}$ is the total natural width of the intermediate autoionizing state, which is the sum of the radiative and autoionization widths: $\Gamma_{d}=A^{d}_{r}+A^{d}_{a}$ (note that rates and corresponding line widths are equivalent in atomic units). Here, it is important to note that the subscripts $r$ and $a$ represent the radiative and autoionization processes, respectively. $L_{d}(E)$ is the Lorentzian line shape function, expressed as
@@
The total rate coefficients for thermal plasmas are described by the following equation, derived by summing across all possible autoionization channels and averaging over the Maxwellian distribution of electron energies [53].
@@
In this expression, $k$ denotes the Boltzmann constant, $h$ is the Planck constant, $T$ represents the electron temperature, and $E$ is the resonance energy.
@@ III Results
The MCDHF method, combined with the relativistic configuration interaction (RCI) approach, is used to calculate the energies of the auto-ionizing and radiative decay states of Fe+, as well as the ground states of Fe2+, as shown in Table 1 and Table 2, respectively. The RCI calculations account for the Breit interaction and approximate QED corrections using the rci code [38]. To assess the impact of QED effects, we performed additional tests and found that, as expected for low-charged ions such as Fe2+, these corrections contribute less than 0.5 meV. Given the present context, their influence on the final results is negligible.
Transition Present Nahar [54] Johansson [55]
and ref. therein
$3d^{6}4s$[6D9/2] $\rightarrow$ $3d^{6}$[5D4] 15.9171 16.0302 16.1353
$3d^{6}$[5D4] $\rightarrow$ $3d^{5}4s4d$[6D1/2] 1.8298 1.9674 1.7633
$3d^{6}$[5D4] $\rightarrow$ $3d^{5}4p^{2}$[6S5/2] 1.7069
$3d^{5}4s4p$[8P5/2] $\rightarrow$ $3d^{6}$[5D4] 9.9064 10.5893 9.6562
$3d^{6}4p$[4F9/2] $\rightarrow$ $3d^{6}$[5D4] 10.5517 10.8410 10.6437
@@
Config. state Present (eV) Ekberg (NIST) [56, 57] (eV)
@@
Since electron correlations play a crucial role in this system, the ASFs are expanded using a large CSF basis. In the present calculation scheme, the CSFs are generated via single and double (SD) excitations of electrons from the Fe2+ ($3d^{6}$ [5D]) valence orbitals to auto-ionizing states of Fe+ ($3d^{5}4s4d$ [6D] and $3d^{5}4p^{2}$ [6S]). A similar approach is applied for radiative decay states, where SD excitations of electrons from the $3d^{6}$ [5D] valence orbitals to the radiative decay states $3d^{5}4s4p$ [8P] and $3d^{6}4p$ [4F] are considered. This method is also used to calculate transition energies from the ground state of Fe+ ($3d^{6}4s$ [6D]) to Fe2+ ($3d^{6}$ [5D]). In these calculations, the spectroscopic orbitals in $3d$, $4s$, and $4p$ were kept open, depending on the states under consideration. This means that, during the generation of CSFs, electrons in these orbitals were allowed to undergo excitations to virtual orbitals as part of the CSF expansion. Valence-valence and core-valence correlations were included, and the convergence of the transition energies and radiative decay rates was monitored through a layer-by-layer expansion. The active space of virtual orbitals was optimized up to the $8k$ subshell. An example of this convergence is shown in Figure 2, where the energies of the $3d^{5}4s4p$ [8P5/2] and $3d^{6}4p$ [4F9/2] states are seen to stabilize around the virtual orbital principal quantum number $n=8$. We compare our result with the theoretical work of Nahar and Pradhan [54] and experimental work of S. Johansson ([55] and references therein), and it is observed that the present values agree well with the experimental value of S. Johansson ([55] and references therein) and demonstrate an improvement upon earlier theoretical work [54].
The number of CSFs for all the considered states is summarized in Table 3 along with the total angular momentum taken in the calculations as well.
State J Number of CSFs
$3d^{6}$ [${}^{5}D$] 00 to $4$ 263,382
$3d^{6}4s$ [${}^{6}D$] $1/2$ to $9/2$ 863,104
$3d^{5}4s4d$ [${}^{6}D$] $1/2$ to $9/2$ 5,714,055
$3d^{5}4p^{2}$ [${}^{6}S$] $5/2$ 957,456
$3d^{5}4s4p$ [${}^{8}P$] $1/2$ to $17/2$ 5,986,145
$3d^{6}4p$ [${}^{4}F$] $1/2$ to $15/2$ 2,955,693
Table 3: Number of CSFs considered for different states in the present calculations.
@@
Figure 2: Calculated energies (in eV) for the states $3d^{5}4s4p$[8P5/2] and $3d^{6}4p$[4F9/2] monitored via layer-by-layer expansion. Black and red lines indicate present and previous studies done on the $3d^{5}4s4p$[8P5/2] and $3d^{6}4p$[4F9/2] states respectively.
@@
The ionization energy of Fe+ ($3d^{6}4s$ [6D9/2]) to Fe2+ ($3d^{6}$ [5D4]) is calculated to be 15.917eV, differing by only 1.3% from the experimental values (see[55] and references therein). The autoionization level energies of the states $3d^{5}4s4d$ [6D1/2] and $3d^{5}4p^{2}$ [6S5/2] from $3d^{6}$ [5D4] are calculated to be 1.830eV and 1.707eV, respectively, agreeing within 3.7% of the experimental energies. Similarly, our calculated radiative decay levels of $3d^{5}4s4p$ [8P5/2] and $3d^{6}4p$ [4F9/2] from $3d^{6}$ [5D4] are 9.9064eV and 10.5517eV, respectively, differing by 2.5% and 0.8% from experimental values. Since our calculations incorporate a more comprehensive set of relevant CSFs, the present results are expected to be more accurate than previous theoretical predictions [54]. These energy levels for various DR states are graphically represented in Fig.3. Table 1 provides a detailed comparison of these energy values with earlier studies. It is evident from the table that, apart from the $3d^{6}4s$ [6D9/2] $\rightarrow$ $3d^{6}$ [5D4] transitions, the present results show better agreement with experimental values compared to prior theoretical data.
@@
The ground-state energies for Fe2+ are also calculated and presented in Table 2. These values are compared with the experimental data provided by Ekberg [56] in the NIST database [57]. The calculated results exhibit excellent agreement with experimental data across all states. As discussed later, these energies were subsequently used in the FAC to generate accurate DR spectra. This underscores the enhanced accuracy and reliability of the current computational approach, demonstrating its capability to predict transitions and align well with experimental observations.
@@
Figure 4: Total DR cross section for Fe2+ recombining into Fe+ is plotted against the relative electron energy. The Lorentzian line shapes (eq.  13) are convoluted with a Gaussian function with a 10-meV width for a better representation of the spectra. The inset presents an enlarged view of the spectrum in the range from 0 to 0.25 eV.
@@
We present the total DR cross section for Fe2+ recombining to Fe+, illustrated in Figure 4. The calculations employ the FAC code to derive the DR spectrum for two autoionizing states of Fe+, namely, $3d^{5}4s4d$ and $3d^{5}4p^{2}$. To enhance the resolution, Lorentzian line shapes, as described in equation 13, are convolved with a Gaussian function characterized by a width of 10 meV. The resulting spectrum exhibits pronounced resonances below 1 eV for both autoionizing states (especially for the $3d^{5}4s4d$ state), with a gradual attenuation of the resonances. Notably, the contributions from the $3d^{5}4s4d$ state are much stronger than from the $3d^{5}4p^{2}$ state. We observe a significant resonance peak at around 0.05 eV, attributed to the $3d^{5}4s4d$ state, along with a couple of resonance peaks between 0.1 to 0.2 eV. Beyond 0.2 eV, the spectrum is almost completely attenuated, with a small resonance peak at 0.4 eV due to the $3d^{5}4p^{2}$ state. There is also a small resonance peak at 0.66 eV, resulting from the contribution of both autoionizing states $3d^{5}4s4d$ and $3d^{5}4p^{2}$. Since the energy levels calculated using the FAC code have an estimated uncertainty of a few eV, they are not highly accurate, necessitating adjustments to improve their precision. To achieve this, we replaced the FAC-calculated energies with more accurate values obtained from large-scale MCDHF calculations using the GRASP code. This substitution ensures more reliable energy positions for low-lying DR resonances. Further details on this energy adjustment procedure can be found in the FAC manual [58]. While this approach enhances the accuracy of resonance positions, some uncertainty remains. In the present case, the ionization energy from Fe II to Fe III is calculated to be 15.9171 eV, which underestimates the experimental value by approximately 0.22 eV. This discrepancy suggests the possibility that the entire DR spectrum may be systematically shifted toward higher energies by a similar amount when compared to experimental data. Given that the deviations between theoretical and experimental values for various autoionizing states fall within this range, it is reasonable to infer that the overall uncertainty in the DR spectrum lies between the present theoretical prediction and a potential upward shift of approximately 0.22 eV. In addition to energy uncertainties, there are inherent uncertainties in the computed rates and cross sections. As noted by Gu in his FAC code [58], for near-neutral ions, such as Fe2+, the uncertainties in rates and cross sections can be as large as 20% or more. Therefore, it is reasonable to assume that the rates and cross sections in the present work have a minimum uncertainty of 20%. The use of the MCDHF method in calculating resonance energies and radiative rates has proven effective in previous studies, yielding results that align well with experimental data [51, 59, 60]. This agreement underscores the validity of the theoretical approach employed in this study. Since a significant amount of data was generated in the present calculation, the energies, resonance strengths, along with their respective widths, are provided as supplementary data.
@@
The total DR rate coefficients for Fe2+ are shown in the figure as a function of temperature on Fig. 5. At high temperatures, the DR rate coefficient decreases according to a $\sim T^{-3/2}$ dependence. At low temperatures, however, the DR process is dominated by a few resonances located just above the ionization limit, which contribute significantly to the rate at low temperatures. This occurs because the $\sim T^{-3/2}$ factor compensates for the low temperature. However, as the temperature gets very low, the exponential cut-off eventually overwhelms the $\sim T^{-3/2}$ increase, causing the rate to decline sharply. As discussed above, the maximum uncertainty in the energy scale is approximately 0.22 eV, which corresponds to an uncertainty of about 2500 K in the temperature scale. Although this may appear significant at first glance, most of the DR rate coefficients contribute predominantly beyond the $10^{4}$ K temperature range [61]. Therefore, this level of uncertainty remains within acceptable limits. Given that the rate coefficients have a minimum uncertainty of 20%, applying the error propagation rule from Eq. II yields a minimum overall uncertainty of approximately 28% in the rate coefficient.